Find an Equation of the line with slope 9 an d containing the point (-1,-5)

Answer by ortarico(23)

Find an Equation of the line with slope 9 an d containing the point (-1,-5)
--------------


(y-y1)=m(x-x1)
y-(-5)=9(x-1)
5y=9x-9
9x-5y-9=0